# Context-Selective State Space Models: Feedback is All You Need

### v1

@@ 3.2 The COFFEE Model
$\Sigma^{(i)}=\begin{cases}x(k)^{(i)}=(I+A^{(i)}\mathrm{diag}(\Delta_{i}(k)))\,x(k-1)^{(i)}+\Delta_{i}(k)\,u(k)_{i},\\ y(k)^{(i)}=C^{(i)}x(k)^{(i)},\end{cases}$ (7)
where the state-dependent gates $\Delta_{i}(k)$ are given in Eq. (5). Here, the learnable parameters are the diagonal elements $\{\lambda_{k}^{(i)}\}_{k=0}^{n-1}$ of the matrix $A^{(i)}$, the output matrix $C^{(i)}\in\mathbb{R}^{1\times n}$, and the feedback vector $w_{D}^{(i)}\in\mathbb{R}^{n}$. To ensure numerical stability during training and inference, the choice of the parameters $\lambda_{k}^{(i)}$ is constrained to the interval [2,0][-2,0], which guarantees that the matrix $I+A^{(i)}$ is stable. A schematic representation of the resulting model is shown in Fig. 1.
Remark 2 (Scalable parallel training).
By construction, $A^{(i)}$ is diagonal, and each component of $\Delta_{i}(k)$ depends only on the corresponding component of $x(k-1)^{(i)}$. As a result, the state dynamics in (7) decomposes into $n$ independent scalar recursions. This in turn implies that the state-transition Jacobian of (7)
$J(k)^{(i)}\;:=\;\frac{\partial x(k)^{(i)}}{\partial x(k-1)^{(i)}}$
is diagonal at every time step. The diagonal structure of the Jacobian drastically reduces both memory and computational requirements. In particular, as shown in [33], trajectory computation can be cast as an optimization problem, where diagonal Jacobians allow efficient parallel solvers with total work and memory scaling that are linear in the sequence length $T$ and the state dimension $n$, i.e., $\mathcal{O}(Tn)$, while the parallel depth in $T$ is only $\mathcal{O}(\log T)$. The overall run-time depends on the number of solver iterations required for convergence, which is typically small in practice. This enables efficient training on parallel hardware, in contrast with cubic and quadratic costs in the state dimension that arise with dense Jacobians.
Remark 3.
(Model with output filtering) Notice that the model in (7), can be modified by adding other filtering capabilities on the output without compromising the possibility to parallelize the architecture. In particular, we can transform the output equation in:
@@
where $\gamma(k)^{(i)}\in\mathbb{R}$ is evaluated as:
@@
with $w_{\gamma}^{(i)}\in\mathbb{R}^{n}$ learnable parameters. As a consequence, the model would have other $nD$ learnable parameters. More generally, the output equation, being static, does not represent a problem for parallelization. Thus, arbitrary non-linearities (if beneficial) can be added.
@@ 4 Experiments
In this section, we present numerical experiments on two primary tasks, the Induction Head (including variants) and MNIST. Implementation details are provided in Appendix G and Appendix E for the Induction Head, and in Appendix H for MNIST.
@@ 4.1 Results on the Induction Head Task
In the following, we present results from a series of experiments designed to solve the Induction Head Task, originally introduced in [34] and adopted as one of the motivating tasks in Mamba [20]. This task was proposed to evaluate in-context learning capabilities of modern deep learning architectures. An explanation of the induction head task is reported in Appendix D. In what follows, the adopted vocabulary is $V=\{1,2,3,4,5,6,7\}$.
Let us start by comparing the performance on the induction head task between the COFFEE model and S6. We test the two models on the IH task with $L_{\text{seq}}=16$, $L_{\text{tri}}=1$ and $L_{\text{tar}}=1$, where $L_{\text{seq}}$, $L_{\text{tri}}$ and $L_{\text{tar}}$ are the lengths of the whole sequence, the trigger and the target, respectively. The results are reported in Table 1.
Table 1: Comparison between S6 and COFFEE on the IH task. Different learning rate were adopted for both S6 and COFFEE, the best ones are reported.
Model
$n$
(state dimension)
$\eta$
(learning rate)
$D$
(embedding dimension)
Loss Accuracy
S6 8 0.003 16 0.852 0.68
COFFEE 8 0.01 16 0.000 0.99+
At this point, the two models have exactly the same number of parameters (i.e. #parameters=784). Our model is able to reach $0.99+$ accuracy in just one epoch, namely with only 5’120’000 training sequences. The S6 performance is obtained by running the training algorithm for 100 epochs, namely 512’000’000 training sequences.
@@
Table 2: COFFEE model performance with reduced state and embedding dimensions
Model
$n$
(state dimension)
$\eta$
(learning rate)
$D$
(embedding dimension)
Loss Accuracy
COFFEE 1 0.01 9 0.000 0.99+
Our model has only 99 learnable parameters. The results are obtained within seven epochs, leading to 35’840’000 training sequences.
To further assess the capabilities of our model, we test it on a variation of the induction head task. The standard implementation of the induction head task, as explained in Appendix D, assumes a sequence with the following structure:
@@
Table 3: COFFEE model performance on this modified version of the IH task
Model
$n$
(state dimension)
$\eta$
(learning rate)
$D$
(embedding dimension)
$L_{noise}$
(noise length)
Loss Accuracy
COFFEE 8 0.003 16 1 0.000 0.99+
COFFEE 8 0.003 16 2 0.002 0.99+
In the above table, “noise length" denotes the length of the noise subsequence between the trigger and target subsequences.
We now test our model on the induction head task with $L_{\text{seq}}=16$, $L_{\text{tri}}=1$ and $L_{\text{tar}}=2$. Results are reported in Table 4.
Table 4: COFFEE model performance with $L_{\text{tri}}=1$ and $L_{\text{tar}}=2$
Model
$n$
(state dimension)
$\eta$
(learning rate)
$D$
(embedding dimension)
Loss Accuracy
COFFEE 8 0.01 16 0.004 0.99
Similarly, with $L_{\text{seq}}=16$, $L_{\text{tri}}=2$ and $L_{\text{tar}}=1$ we obtain the results in Table 5.
Table 5: COFFEE model performance with $L_{\text{tri}}=2$ and $L_{\text{tar}}=1$
Model
$n$
(state dimension)
$\eta$
(learning rate)
$D$
(embedding dimension)
Loss Accuracy
COFFEE 8 0.003 16 0.660 0.78
To conclude, we test our model on the induction head task with $L_{\text{seq}}=32,64,128,256$, $L_{\text{tri}}=1$ and $L_{\text{tar}}=1$. The results and settings are reported in Table 6.
Table 6: COFFEE model performance with increased sequence length
Model
$L_{\text{seq}}$
(sequence length)
$n$
(state dimension)
$\eta$
(learning rate)
$D$
(embedding dimension)
Loss Accuracy
COFFEE 32 8 0.01 16 0.000 0.99+
COFFEE 64 8 0.01 16 0.001 0.99+
COFFEE 128 8 0.01 16 0.000 0.99+
COFFEE 256 8 0.01 16 0.001 0.99+
Considering that S6 is the first SSM to introduce selectivity and is regarded as the state of the art, we conclude that our model deserves attention and further investigation. Based on the results above, COFFEE appears to be significantly more effective compared to the S6 model.
@@ 4.2 Results on the MNIST task
To further assess our model ability to capture dependencies between tokens, we tested our SSM on the MNIST dataset. The adopted architecture and the hyperparameters setting are fully explained in Appendix H. In Table 7 we report a summary of the results obtained: in the table, “COFFEE with output filtering” refers to the usual COFFEE model improved with filtering capabilities on the output (see Remark 3).
Table 7: Performance comparison between COFFEE and S6 using the MNIST dataset
Model Epochs
$n$
(state dimension)
# parameters Loss Accuracy
S6 100 2 5585 1.891 28.2%
S6 100 16 10085 1.876 28.6%
COFFEE 100 2 3385 0.107 96.6%
COFFEE with output filtering 100 2 3585 0.100 97.0%
@@
This experiment demonstrates the practical applicability of COFFEE in a scenario that is very different from those considered in the previous section. As discussed in [20], selectivity is a fundamental property of powerful architectures for sequential modeling. In the previous section we established the relevance and efficiency of the selection mechanism introduced in COFFEE for tasks in the induction head family; with the results summarized in Table 7, we show how this embedded mechanism transfers to a completely different and more complex task, without incurring a prohibitive increase in model size.
@@
In summary, we demonstrate that our model is flexible and able to outperform the current S6 model in significantly different tasks: for these reasons, we believe it deserves consideration as a building block in more complex architectures, such as Mamba [20].
5 Insights on the functioning principles of COFFEE
One of the major challenges in deep learning is related to the mechanistic interpretability of the learned models, the latter being essential towards the safe and responsible application of these tools in critical domains. In order to understand how these models encode and process information, the first crucial step is to gain awareness of the internal mechanisms and limitations of trained mathematical models.
@@ 5 Insights on the functioning principles of COFFEE
Our COFFEE model has been designed with a few basic, interpretable functioning principles in mind:
  • P1:
@@
    The state represents the memory of the system, which is propagated to the next iteration by the state matrix.
  • P2:
@@
    The update gates $\Delta_{i}(k)$ select if and how much of the input $u(k)_{i}$ is transferred to the state dynamics and hence to memory.
  • P3:
@@
    Different areas of the state space allow the next encoded symbol to affect its future evolution differently (context selectivity).
Next, we investigate whether the proposed model is indeed capable of operating by exploiting the previous rules. An in-depth analysis, aimed at identifying the key principle at play, is possible only by considering small models with limited degrees of freedom and simple tasks. Even the simple architecture that solved the inductive head task, with only ninety-nine parameters (see Section 4), turned out to be impractical to study due to the large dimensionality of the space and the lack of convenient visualization of the state and model parameter dynamics. A further recurring difficulty is presented by the fact that a model that is general enough to solve a wide spectrum of tasks can, in general, solve a given problem in many different ways, using inequivalent emerging “strategies”.
@@
We therefore decided to reduce the “complexity” of the induction head task as much as possible, so that we could geometrically represent the evolution of the state and more easily identify the choices of parameters that can implement the principles described above. In this way, we can verify that the learned solution is indeed behaving as expected. The simplified inductive head task (IH0, described in detail in Section F.1) considers sequences of four symbols chosen in $V=\{1,2,3\},$ where $1$ represents the trigger. The trigger appears two times in each sequence, one of which as the last symbol. To solve this problem, we chose embeddings in $\mathbb{R}^{2},$ and a COFFEE model composed of two one-dimensional SSM, so that the evolution can be depicted in $\mathbb{R}^{2}$ as well.
In order to highlight the role of the functioning principles P1-3 above, we introduce further model simplifications:
  • Principle P1 is implemented by choosing the state matrix to be the identity (i.e. all $\lambda^{(i)}=0$ in Equation 7.), thus implementing a perfect memory. Furthermore, we also assume $C^{(i)}=1$ for $i=1,2$, so the state values correspond to the outputs, and simplify interpretation.
  • To implement P2, the update gates $\Delta_{i}(k)$ are chosen as functions of the corresponding state features: Given their definition in (5), the sigmoid functions play a key role. By choosing $w_{D}^{(i)}=1$ in our reduced model, the state value remains the only relevant parameter. High positive values of the state components correspond to update gates close to one, and negative values bring the update gates close to zero.
  • Regarding P3, in the simplified model the state belongs to $\mathbb{R}^{2},$ and its components directly activate the corresponding sigmoids. Loosely speaking, a state in the upper half-plane will cause the next input to “save” its second component (feature) in memory, and the right half-plane will memorize the first one. The first quadrant will save both, and the third quadrant neither.
@@
With these simplifications in place, the only parameters left to learn are the embeddings. Taking into account the previous observations and the fact that the inductive head task consists of memorizing the symbol following the first trigger, we can intuitively guess the general form of a successful embedding. In particular, the trigger symbol embedding shall move the state in a region of activation of the sigmoid that is high and neutral with respect to the embedding of the target symbols. Natural candidates are vectors on the bisector of the first quadrant. The embeddings of the target symbols shall (i) be well distanced, in order to ensure distinguishability of the output; (ii) move the state from the triggered position to an area that is closer to their position than the others, so that their symbol is chosen as the final output; (iii) lower the activation of the sigmoid, and hence the update gates, so that further inputs will not cause the state to move too much towards the embedding of different symbols. Natural candidates for the simplified problems are large vectors in the third quadrant.
@@
We would like to verify that the learned solutions for the IH task respect these principles. Once all the proposed model simplifications are implemented, the training phase does not always end successfully due to the extremely limited number of parameters to be optimized (6, corresponding to the embedding), and a suitable initialization must be considered. With this in mind, we trained the simplified model starting from a tentative embedding that satisfies the design principles sketched above, where:
$1$ $\rightarrow$ $[6,6]^{\top}$
$2$ $\rightarrow$ $[-10,-1]^{\top}$
$3$ $\rightarrow$ $[-1,-10]^{\top}.$
@@
From this initial condition, the learning phase converges extremely quickly, and the learned model solves the task with $100\%$ accuracy. The learned solution corresponds to the embedding:
$1$ $\rightarrow$ $[5.394,5.3433]^{\top}$
$2$ $\rightarrow$ $[-10.2643,-1.5753]^{\top}$
$3$ $\rightarrow$ $[-1.5395,-10.3404]^{\top}.$
The full trajectories of the state variables, as they react to input and context, are depicted in Figure 2 in the Appendix. Their behavior is as expected, in accordance with principles P1-3.
In light of this, the question “Is the automatically learned solution exploiting the same mechanism and degrees of freedom we describe in P1-2-3?" can be answered positively, at least in the case of the simplified model composed of perfect memories (integrators). In the general case, the introduction of nontrivial state dynamics induces a drift field which further complicates the analysis, and feedback matrices $w_{D}$ influence nontrivially the geometry of the activation areas, and yet a general agreement with the proposed functioning principles can be seen, albeit not as clearly.

### v2

@@ 3.2 The COFFEE Model
$\begin{cases}x(k)^{(i)}=(I+A^{(i)}\mathrm{diag}(\Delta_{i}(k)))\,x(k-1)^{(i)}+\Delta_{i}(k)\,u(k)_{i},\\ y(k)^{(i)}=C^{(i)}x(k)^{(i)},\end{cases}$ (7)
where the state-dependent gates $\Delta_{i}(k)$ are given in (5). Here, the learnable parameters are the diagonal elements $\{\lambda_{k}^{(i)}\}_{k=0}^{n-1}$ of the matrix $A^{(i)}$, the output matrix $C^{(i)}\in\mathbb{R}^{1\times n}$, and the feedback vector $w_{D}^{(i)}\in\mathbb{R}^{n}$. To ensure numerical stability during training and inference, we enforce stability for the matrix $I+A^{(i)}$. A schematic representation of COFFEE is shown in Fig. 1.
Remark 2 (Model with output filtering).
Notice that the model in (7), can be extended by introducing additional filtering mechanisms at the output level. In particular, we can transform the output equation in:
@@
where $\gamma(k)^{(i)}\in\mathbb{R}$ is given by:
@@
and $w_{\gamma}^{(i)}\in\mathbb{R}^{n}$ are learnable parameters. This extension introduces an additional $nD$ learnable parameters.
Notably, since the output equation is static, this generalization does not compromise the possibility, discussed in the next section, of parallelizing the model. This remains true even if other nonlinearities are added to the output equation.
4 Parallelization
Architectural parallelization is a fundamental requirement to fully exploit the computational capabilities of modern GPUs, including Tensor and CUDA cores. However, the model described in Equation (7) does not naturally admit an efficient parallel implementation. Indeed, computing the state $x(k)$ requires access to the previous state x(k1)x(k-1), thus pointing to an intrinsic sequential dependency.
Moreover, when implemented using standard automatic differentiation frameworks (e.g., PyTorch, JAX), this approach leads to the construction of a large computational graph. Such a graph not only incurs significant memory overhead but may also cause vanishing gradients or numerical instabilities during gradient backpropagation.
To overcome these limitations, we adopt a parallelization strategy originally introduced by [27] and subsequently refined by [13]. In [27], an algorithm known as DEER is developed to address both the discrete-time and continuous-time formulations. When specialized to the discrete-time recurrent neural network setting, such framework allows inference to be reformulated as the solution of a fixed-point problem, which can then be efficiently solved using Newton’s method. To use this theory in our setting, we write the state equation of (7) in the form
$x(k)=f_{k}(x(k-1))$
where $x(k)\in\mathbb{R}^{\tilde{n}}$ may denote each single component $x(k)^{(i)}$ (in which case $\tilde{n}=n$), or the whole state obtained by stacking the $x(k)^{(i)}$ on top of each others (in which case $\tilde{n}=Dn$); in fact, depending on the hardware and on the values of $n$ and $D$, it may be convenient to consider each $\Sigma^{(i)}$ separately or to group them together as a unique SSM. The function $f_{k}$ represents the nonlinear state transition function and its dependence on $k$ also accounts for the contribution of $u(k)$. Given a collection of candidate states $\{x(k)\}_{k=1}^{L}$, where $L$ denotes the sequence length, we define the residual as:
$r(x_{1:L})\;:=\;[x(1)-f_{1}(x(0)),x(2)-f_{2}(x(1)),...,x(L)-f_{L}(x(L-1))]$
[…]
Building on the previous considerations, our model has been designed so that the Jacobians are inherently diagonal, making quasi-DEER an exact implementation rather than an approximation of DEER. Specifically, in our model, $A^{(i)}$ is diagonal, and each component of $\Delta_{i}(k)$ depends solely on the corresponding component of $x(k-1)^{(i)}$. Consequently, the state dynamics in (7) decomposes into $\tilde{n}$ independent scalar recursions. This further implies that the state-transition Jacobian is diagonal at every time step.
@@ 5 Experiments
We report numerical experiments on two primary tasks: Induction Head (including variants) and MNIST. Implementation details are provided in Appendix H and Appendix F for the Induction Head, and in Appendix I for MNIST.
@@ 5.1 Results on the Induction Head task
In the following, we present results from a series of experiments designed to solve the Induction Head (IH) Task, originally introduced in [33] and adopted as one of the motivating tasks in Mamba [15]. This task was proposed to evaluate in-context learning capabilities of modern deep learning architectures. An explanation of the induction head task is reported in Appendix E. In what follows, the adopted vocabulary is $V=\{1,2,3,4,5,6,7\}$.
We first compare the performance of COFFEE and S6 on the IH task with $L_{\text{seq}}=16$, $L_{\text{tri}}=1$ and $L_{\text{tar}}=1$, where $L_{\text{seq}}$, $L_{\text{tri}}$ and $L_{\text{tar}}$ are the lengths of the whole sequence, the trigger and the target, respectively. As shown by the results in Table 1, COFFEE is able to reach $0.99+$ accuracy in just one epoch, namely with only 5’120’000 training sequences. The S6 performance is obtained by running the training algorithm for 100 epochs (512’000’000 training sequences).
Table 1: Comparison between S6 and COFFEE on the IH task. Different learning rate $\eta$ were adopted for both S6 and COFFEE, the best ones are reported. For both models the state dimension is $n=8$ and the embedding dimension is $D=16$.
Model $\eta$ # par. Epochs Loss Acc.
S6 0.003 784 100 0.852 0.68
COFFEE 0.01 512 1 0.000 0.99+
@@
Table 2: COFFEE performance on the IH task with reduced state and embedding dimensions. The learning rate is $\eta=0.01$. The embedding dimension is denoted by $D$, the state dimension by $n$ and # par. denotes the number of learnable parameters.
Model $n$ $D$ # par. Loss Accuracy
COFFEE 1 8 96 0.004 0.99+
COFFEE 1 4 48 0.002 0.99+
COFFEE 1 2 24 0.035 0.99
COFFEE 1 1 12 1.793 0.16
We see that COFFEE achieves near perfect accuracy with only 24 parameters which is a remarkable performance.
To further assess the capabilities of COFFEE, we test it on a variation of the IH task. The standard implementation of the IH task, as explained in Appendix E, assumes a sequence with the following structure:
@@
Table 3: COFFEE model performance on a variation of the IH task. Here the state dimension is $n=8$, the embedding dimension is $D=16$ and the learning rate is $\eta=0.003$. The length of the noise subsequence between the trigger and target subsequences is denoted by $L_{\text{noise}}$.
Model $L_{\text{noise}}$ Loss Accuracy
COFFEE 1 0.000 0.99+
COFFEE 2 0.002 0.99+
We also evaluate COFFEE on the IH task under different trigger and target lengths: $L_{\text{tri}}=1$, $L_{\text{tar}}=2$ and $L_{\text{tri}}=2$, $L_{\text{tar}}=1$ while keeping $L_{\text{seq}}=16$. Results are in Table 4.
Table 4: COFFEE model performance with different values of $L_{\text{tri}}$ and $L_{\text{tar}}$. Here the state dimension is $n=8$, the embedding dimension is $D=16$ and the learning rate is denoted by $\eta$.
Model $\eta$ $L_{\text{tri}}$ $L_{\text{tar}}$ Loss Accuracy
COFFEE 0.01 1 2 0.004 0.99
COFFEE 0.003 2 1 0.660 0.78
To conclude, we test our model on the IH task with increasing sequence lengths $L_{\text{seq}}=32,64,128,256$, $L_{\text{tri}}=1$ and $L_{\text{tar}}=1$. The results and settings are reported in Table 5.
Table 5: COFFEE model performance with increased sequence length. Here the state dimension is $n=8$, the embedding dimension is $D=16$ and the learning rate is $\eta=0.01$. The length of the sequence is denoted by $L_{\text{seq}}$.
Model $L_{\text{seq}}$ Loss Accuracy
COFFEE 32 0.000 0.99+
COFFEE 64 0.001 0.99+
COFFEE 128 0.000 0.99+
COFFEE 256 0.001 0.99+
Overall, the experimental results show that COFFEE outperforms S6 across multiple IH settings, attaining higher accuracy while requiring significantly fewer parameters and training sequences. Given that S6 represents a strong baseline among selective SSMs, these findings highlight the potential of state-feedback mechanisms for improving performance on in-context learning tasks.
Ablation studies. COFFEE was derived through three main steps. First, we linearized S6; second, we applied a state-feedback; third, we performed a change of basis to eliminate redundancy. A natural question is then which of these three steps contributes the most to the overall performance. To address this question, Table 6 reports experimental results on the IH task with $L_{\text{seq}}=16$, $L_{\text{tri}}=1$, and $L_{\text{tar}}=1$ considering different “partial" models.
Table 6: Performance of the S6 model, COFFEE with linearization only as in Step 1, COFFEE obtained by applying Steps 1 and 3 but no state feedback (COFFEE no fb.), and the full model (COFFEE final.). To assess the importance of the nonlinearity in the state feedback (the sigmoid in (5)), we also tested the model without it; the corresponding results are reported as COFFEE lin. fb..
Model $n$ $D$ # par. Loss Accuracy
S6 8 16 784 0.852 0.69
COFFEE as in (4) 8 16 784 0.906 0.68
COFFEE no fb. 8 16 512 0.788 0.74
COFFEE final (7) 8 16 512 0.000 0.99+
COFFEE lin. fb. 8 16 512 1.845 0.16
Remark 3.
All experiments reported in this section were repeated multiple times, with negligible standard deviation.
@@ 5.2 Results on the MNIST task
To further assess our model ability to capture dependencies between tokens, we tested our SSM on the MNIST dataset. For each model, we employ a single state-space layer without feedforward, with the same number of features. The adopted architecture and the hyperparameters setting are fully explained in Appendix I. In Table 7 we report a summary of the results obtained: in the table, “COFFEE + OF” refers to the usual COFFEE model improved with filtering capabilities on the output (see Remark 2).
Table 7: Performance comparison between COFFEE and S6 using the MNIST dataset. Here, we consider $100$ Epochs. We denote by $n$ the state dimension and by # par. the overall number of learnable parameters of the model. Experiments were repeated ten times.
Model $n$ # par. Loss Acc. Std
S6 2 5585 1.891 28.2% 0.1
S6 16 10085 1.876 28.6% 0.1
COFFEE 2 3385 0.107 96.6% 0.1
COFFEE + OF 2 3585 0.100 97.0% 0.1
@@
This experiment demonstrates the practical applicability of COFFEE in a scenario that is very different from those considered in the previous section, and show how the S6 performance is affected once the feedforward layer is removed. As discussed in [15], selectivity is a fundamental property of successful architectures for sequential modeling. In the previous section we established the relevance and efficiency of the selection mechanism introduced in COFFEE for tasks in the induction head family; with the results summarized in Table 7, we show how this embedded mechanism transfers to a completely different and more complex task, without incurring a prohibitive increase in model size.
@@
In summary, we demonstrate that COFFEE is flexible and outperforms S6 in very different tasks: for these reasons, we believe it deserves consideration as a building block in complex architectures, such as Mamba [15].
6 Functioning principles of COFFEE
@@
The COFFEE model has been designed with a few basic, interpretable functioning principles in mind:
@@
P1: The state represents the memory of the system, which is propagated to the next iteration by the state matrix.
@@
P2: The update gates $\Delta_{i}(k)$ select if and how much of the input $u(k)_{i}$ is memorized in the state dynamics.
@@
P3: Different areas of the state space allow the next encoded symbol to affect its future evolution differently (context selectivity).
@@
In order to investigate whether the proposed model is indeed operating by exploiting these principles, we reduced the “complexity” of the induction head task as much as possible, so that we could geometrically represent the evolution of the state. In fact, even the simple architecture that solved the inductive head task with 24 parameters (see Section 5), turned out to be impractical to study due to the lack of convenient visualization of the parameter dynamics. The simplified inductive head task (IH0, described in detail in Appendix G.1) considers sequences of four symbols chosen in $V=\{1,2,3\},$ where $1$ represents the trigger. The trigger appears two times in each sequence, one of which as the last symbol. To solve this problem, we chose embeddings in $\mathbb{R}^{2},$ and a COFFEE model composed of two one-dimensional SSM, so that the evolution can be depicted in $\mathbb{R}^{2}$ as well.
In order to highlight the role of the functioning principles described above, we introduce further model simplifications: For P1, we choose all $A^{(i)}=0$ in Equation (7)), thus implementing a perfect memory. Furthermore, we also assume $C^{(i)}=1$ for $i=1,2$, so that the state values correspond to the outputs, and simplify interpretation. For P2, we impose $w_{D}^{(i)}=1$, so that the state value remains the only relevant parameter. High positive values of the state components correspond to update gates close to one, and negative values bring the update gates close to zero. For P3, a state in the upper half-plane will cause the next input to “save” its second component (feature) in memory, and the right half-plane will memorize the first one. The first quadrant will save both, and the third quadrant neither. With these restrictions in place, the only parameters left to learn are the embeddings.
@@
We can now intuitively guess the general form of a successful embedding: In particular, the trigger symbol embedding shall move the state in a region of activation of the sigmoid that is high and “neutral” with respect to the embedding of the target symbols. Natural candidates are vectors on the bisector of the first quadrant. The embeddings of the target symbols shall (i) be well distanced, in order to ensure distinguishability of the output; (ii) move the state from the triggered position to an area that is closer to their position than the others, so that their symbol is chosen as the final output; (iii) lower the activation of the sigmoid, and hence the update gates, so that further inputs will not cause the state to move too much towards the embedding of different symbols. Natural candidates are then large vectors in the third quadrant.
@@
We next verify that the learned embeddings for the IH task respect these principles. We trained the simplified model starting from a tentative embedding that satisfies the design principles sketched above, where:
$\vskip-5.69054pt1\rightarrow[6,6]^{\top},\quad 2\rightarrow[-10,-1]^{\top},\quad 3\rightarrow[-1,-10]^{\top}.$
@@
From this initial condition, which does not always yield the correct results, the learning phase converges quickly, and the learned model solves the task with $100\%$ accuracy. The learned solution corresponds to the embedding:
$1\!\rightarrow[5.4,5.3]^{\top},\,2\!\rightarrow[-10.4,-1.6]^{\top},3\!\rightarrow[-1.5,-10.3]^{\top}$
The full trajectories of the state variables are depicted in Figure 2 in the Appendix. Their behavior shows the design principles P1-3 in action. It is worth noting that in this simple example, despite the existence of such optimal solution, the training phase does not always end successfully due to the extremely limited number of parameters to be optimized (6, corresponding to the embedding), and considering a suitable initialization is key [^5] These issues do not emerge for the fully-parametrized model.. Increasing the number of parameters quickly improves the training success, providing clear evidence of the beneficial effects of introducing redundancies in number of trainable parameters, despite being inefficient in terms of resources. In this sense, P1-P3 represent both a way to interpret the behavior of the model, and a useful starting point to propose effective initializations, even in the presence of fewer redundant parameters.